# Color it, Code it, Cancel it: $k$-local dynamical decoupling from classical additive codes

### v1

@@ V.2.1 Linear projective geometry code
A projective geometry space $\rm PG(n,q)$ is the set of equivalence classes of the vector space $\mathbb{F}_{q}^{n+1}/\{\bm{0}\}$ in which two vectors $\bf{u},\bf{v}\in\mathbb{F}_{q}^{n+1}$ belong to the same equivalence class $\bf{u}]$ if they differ by scalar multiplication, i.e., $\bf{u}=\alpha\bf{v}$ for some $\alpha\in\mathbb{F}_{q}$. Consequently, one can view each point $\bf{u}]\in\rm PG(n,q)$ as a collection of points in $\mathbb{F}_{q}^{n+1}$ starting from $\bf{u}$ and passing through the origin.
The projective space that we will use to construct the decoupling group is $\rm PG(n,4)$. The decoupling group is constructed using the check matrix $H$ of the linear code $\mathcal{C}$. As a reminder, the codeword $\bf{c}\in\mathcal{C}$ is the kernel of the check matrix, i.e.
@@
As a first step, we show how to construct a DD group for universal cancellation of all two-body interactions from the projective space. We begin by choosing the columns of the check matrix $H$ to correspond to points $\bf{u}]$ in the projective space $\rm PG(n,4)$; that is, to equivalence classes of nonzero vectors in $\mathbb{F}_{4}^{n+1}$. The total number of such equivalence classes is given by
@@
Because the check matrix $H$ of the linear code $\mathcal{C}$ is the generator matrix for the dual code $\mathcal{C}^{\perp}$, using Theorem IV.1, we can construct an OA from $\mathcal{C}^{\perp}$ with parameters
@@
An important remaining question is how to construct a cap set in $\rm PG(n,4)$ with the largest possible size. In general, this is a challenging problem in projective geometry. However, explicit constructions of maximal cap sets are known for small values of $n$ [72]. For example, the largest cap set sizes for $n=2,3,4$ are $n_{H}=6,17,41$ respectively. For higher dimensions $n$, near-optimal constructions are known, although it is unlikely that physically relevant graphs of solid-state devices will require these large chromatic number $\chi_{I}$.
@@ V.2.2 Additive projective geometry code
In the preceding construction, we employed linear projective geometry codes to design universal DD groups for two-local and three-local interactions. We now show that if we use additive projective geometry codes [73], we can construct more time-efficient sequences. The key insight is that the field $\mathbb{F}_{4}$, when considered under addition, is isomorphic to $\mathbb{F}_{2}\times\mathbb{F}_{2}$. Exploiting this structure, Ref. [73] demonstrates that the check matrix $H$ of an additive code with parameters $(\chi_{I},2^{k},d)_{4}$ can be constructed from a collection of $\chi_{I}$ lines in the projective space $\rm PG(2\chi_{I}-k-1,2)$, such that any d1d-1 lines in the set are linearly independent.
To universally decouple two-local interactions, we require the code $\mathcal{C}$ to have distance $d=3$. Consequently, we need to choose a collection of $\chi_{I}$ lines in the projective space $\rm PG(2\chi_{I}-k-1,2)$ such that any $2$ lines in the set are linearly independent. This set is known in projective geometry as (partial) 1-spread. The maximum logical dimension $k$ is given by
@@
where $h$ is the smallest integer number such that $2^{h}\geq 3\chi_{I}+1$ for even $h$, and $2^{h}\geq 3\chi_{I}+5$ for odd $h$. As a result, the orthogonal array constructed from $\mathcal{C}^{\perp}$ has parameter
$\text{OA}(2^{\min(\lceil\log_{2}(3\chi_{I}+1)\rceil,\lceil\log_{2}(3\chi_{I}+5)\rceil)},\chi_{I},4,2).$ (52)
As a simple example, Table 2(a) lists a full 1-spread of the projective space $\rm PG(3,2)$, where each projective line is represented by two projective points. Each line is further associated with a colored qubit $c_{i}$. In Table 2(b), we map each projective line in $\rm PG(3,2)$ to a projective point in $\rm PG(3,4)$ via a simple embedding from $\mathbb{F}_{2}\times\mathbb{F}_{2}$ to $\mathbb{F}_{4}$
@@
The resulting rows in the table form generators of a DD group capable of suppressing arbitrary two-local interactions.
We note that the above construction uses the same number of colored qubits and time slices in the decoupling sequence as the multi-axis protocol introduced in Ref. [27]. This is not a coincidence. The multi-axis protocol can be understood as implicitly employing an additive code over $\mathbb{F}_{4}$. This connection arises because the Hadamard matrix defines a binary linear code (the Hadamard code), whose rows or columns correspond to codewords. By grouping these rows into Schur sets [74] and mapping from $\mathbb{F}_{2}\times\mathbb{F}_{2}$ to $\mathbb{F}_{4}$, one preserves the additive structure of the binary code. Consequently, our framework naturally includes the protocol from Ref. [27], providing an alternative way to understand its underlying structure.
@@
Figure 7: Scaling of decoupling sequences. Scaling of sequence length $L$ with the chromatic number $\chi_{I}$ for the code constructions listed in Table 3. (a) For comparison, the linear baseline $L=4\chi_{I}$ is shown in black. All codes with distance $d=3$ exhibit linear scaling in $\chi_{I}$, while those with $d=4$ show quadratic scaling, consistent with the scaling reported in Ref. [30].
@@ V.2.3 Tailored sequences for spin-qubit
A similar strategy also applies to additive projective geometry codes: for each line in the projective space, there exist three distinct representations that also avoid Heisenberg exchange terms as codewords. As an illustration, Table 2(c) presents a tailored DD group designed to detect Heisenberg exchange interactions, derived from the construction in Table 2(b). Notably, this construction increases the number of color classes from 5 to 15 while retaining the same number of generators. In the case of the multi-axis protocol introduced in Ref. [27], this modification is equivalent to cyclically permuting the rows in the Schur set.
@@
Let us apply this idea to the quotient graph in Fig. 6(b). We begin by constructing a cap set in $\rm PG(2,4)$, which corresponds to a universal DD sequence that decouples all three-local interactions among six qubits. This cap set is given by the additive generators of the self-dual hexacode [75]
@@
The resulting DD sequence suppresses both the Heisenberg exchange interactions and the scalar chirality terms. However, it stills contains six generators. By mapping the generator set to a minimum covering problem, as described in Sec. IV.4, we find that only four of these generators are necessary to detect all relevant interaction term
@@
We emphasize that the expansion approach is effective primarily for bang-bang sequences. Under bounded control, finite pulse durations can lead to Heisenberg exchange or scalar chirality interactions transforming into other three-local terms that the sequence may not cancel. In such cases, alternative constructions may be required to obtain tailored, time-efficient bounded-control DD sequences. Alternatively, one may resort to the universal DD sequences discussed in Sec. V.2.1, which, although potentially longer, suppress all three-local interactions. Finally, we note that shorter sequences can often be obtained by systematically removing generators from a universal sequence, one at a time, as demonstrated above. This process might reduces the number of supported color classes but preserving the required decoupling properties. Developing such compact, tailored sequences remains an interesting direction for future work.
@@
We also analyze the sequence lengths as a function of the chromatic number $\chi_{I}$. To visualize this, in Fig. 7(a) we show the scaling of different DD sequences that suppress two-local interactions as the chromatic number $\chi_{I}$ increases; in Fig. 7(b), we focus on sequences targeting three-local interactions. In both cases, we compare the scaling behavior of the sequence lengths against ideal linear and quadratic scaling.
For two-local interactions, we find that the universal constructions scale linearly with sequence length $4\chi_{I}$, while the tailored constructions scale more compactly with $2\chi_{I}$. In the case of three-local interactions, both universal and tailored constructions exhibit quadratic scaling $L\propto\chi_{I}^{2}$, but the tailored sequences have a smaller prefactor, leading to more efficient implementations. These scaling behaviors are consistent with those of Ref. [30].
@@ V.3 Digital-analog simulation
In the previous section, we showed how to construct DD sequences that suppress residual interactions within the system and its environment. In this section, we demonstrate how such sequences can also be leveraged for Hamiltonian simulation and digital-analog quantum computing. For concreteness, we consider the Kitaev’s honeycomb lattice toy model, described by the Hamiltonian [76]
@@
where the sum is taken over the red (r), blue (b), and green (g) colored edges in Fig. 8(a). This model is exactly solvable by using a mapping to Majorana fermions and exhibits both gapped and gapless phases depending on whether the coupling strengths $J_{x},J_{y},J_{z}$ satisfy the triangle inequality [76]. Here, we focus on simulating the system deep in the gapless phase, where $J_{x}=J_{y}=J_{z}=J$.
@@
Figure 8: Dynamical quantum simulation of the Kitaev model. (a) A finite-size Kitaev honeycomb lattice, where edges are colored red, blue, and green to represent $X_{i}X_{j}$, $Y_{i}Y_{j}$, and $Z_{i}Z_{j}$ interactions, respectively. The labels 11 through 66 indicate the equivalence classes of qubits under the vertex color partition which respects the edges coloring, not individual physical qubits. (b) The folded Kitaev lattice (i.e., the quotient graph) obtained by collapsing all qubits within the same equivalence class. Compared to the standard unit cell, the quotient graph includes three additional edges (2-5), (3-4), and (1-6) that arise from the folding.
As in previous sections, we begin by identifying the effective number of qubits in Kitaev’s honeycomb model. If we were to universally decouple all interactions, the honeycomb lattice would be trivially 2-colorable. However, because we aim to selectively preserve certain interactions, the effective model becomes more complex.
In particular, due to the link-dependent interaction terms in the Hamiltonian, the symmetry of the model reduces the problem to an effective system involving six qubits. In Fig. 8(a), we color the qubits that are equivalent under the translation symmetry of the honeycomb lattice. The resulting quotient graph, formed by identifying these symmetry-equivalent sites, contains six effective colored qubits. There are three extra edges appear between opposite effective vertices, as shown in Fig. 8(b).
To demonstrate the power of the effective graph in Fig. 8(b), we use the simple algorithm presented in Section IV.4 to find the minimal DD group for the bang-bang sequence. We also explicitly construct a DD group for the bounded sequence, see Appendix D. Using the same terminology as before, we associate $H_{S}^{\parallel}=H_{\rm comb}$. Consequently, the remaining interactions belong to the list $H_{S}^{\perp}$.
@@
$W_{3}=X_{1}Z_{2}Z_{5}X_{6},\quad W_{4}=Y_{2}X_{3}X_{5}Y_{6},$ (61b)
and are closely related to a known symmetry of the Kitaev honeycomb model, namely the plaquette operator [76]
$W_{p}=Z_{1}X_{2}Y_{3}Y_{4}X_{5}Z_{6}.$ (62)
One can verify that $W_{1}W_{2}=W_{p}$ and $W_{3}W_{4}=W_{p}$, that is, each pair provides a decomposition of the original plaquette symmetry. The emergence of these subsymmetries is expected: since the effective Kitaev honeycomb model arises from folding the original lattice, it can exhibit additional symmetries not present in the unfolded geometry.
If we take $W_{p}$ as a generator of a DD group $\mathcal{G}$, the resulting sequence effectively removes the interaction terms along the outer edges of the hexagon in Fig. 8(b), aligning the system’s dynamics with that of the Kitaev model. However, the inner-edge interactions commute with $W_{p}$, so they are not eliminated by this sequence and remain in the effective Hamiltonian. This again reflects the presence of the subsymmetries introduced by lattice folding.
Either by numerically solving the associated minimum-covering problem or by manually verifying it, we find that using either the pair $(W_{1},W_{2})$ or $(W_{3},W_{4})$ as generators is sufficient to construct a DD group $\mathcal{G}$ for the bang-bang sequence. Consequently, the effective Hamiltonian have the same interaction as shown in the edges of Fig. 8(b). However, the global site is opposite of what it should be, i.e $H_{S}^{\parallel}=-H_{\rm comb}$.

### v2

@@ V.2.1 Linear projective geometry code
A projective geometry space $\rm PG(n,q)$ is the set of equivalence classes in the vector space $\mathbb{F}_{q}^{n+1}/\{\boldsymbol{0}\}$ of which two vectors $\bf{u},\bf{v}\in\mathbb{F}_{q}^{n+1}$ belong to the same equivalence class $\bf{u}]$ if they differ by scalar multiplication, i.e. $\bf{u}=\alpha\bf{v}$ for some $\alpha\in\mathbb{F}_{q}$. Consequently, one can view each point $\bf{u}]\in\rm PG(n,q)$ as a collection of points in $\mathbb{F}_{q}^{n+1}$ starting from $\bf{u}$ and passing through the origin.
The projective space that we will use to construct the decoupling group is $\rm PG(n,4)$. The decoupling group is constructed using the linear check matrix $H$ of the linear code $\mathcal{C}$. As a reminder, the codeword $\bf{c}\in\mathcal{C}$ is the kernel of the check matrix, i.e.
@@
As a first step, we show how to construct a decoupling group for universal cancellation of all two-body interactions from the projective space. We begin by choosing the columns of the check matrix $H$ to correspond to points $\bf{u}]$ in the projective space $\rm PG(n,4)$; that is, to equivalence classes of nonzero vectors in $\mathbb{F}_{4}^{n+1}$. The total number of such equivalence classes is given by
@@
Because the check matrix $H$ of the linear code $\mathcal{C}$ is the linear (not additive) generator matrix for the dual code $\mathcal{C}^{\perp}$, using Theorem IV.1, we can construct an OA from $\mathcal{C}^{\perp}$ with parameters
@@
An important remaining question is how to construct a cap set in $\rm PG(n,4)$ with the largest possible size. In general, this is a challenging problem in projective geometry. However, explicit constructions of maximal cap sets are known for small values of $n$ [FU2015]. For example, the largest cap set sizes for $n=2,3,4$ are $n_{H}=6,17,41$ respectively. For higher dimensions $n$, near-optimal constructions are known, although it is unlikely that physically relevant graphs of solid-state devices will require these large chromatic number $\chi_{I}$.
@@ V.2.2 Additive projective geometry code
In the preceding construction, we employed linear projective geometry codes to design universal DD groups for two-local and three-local interactions. We now show that if we use additive projective geometry codes [BLOKHUIS2004161], we can construct more time-efficient sequences. The key insight is that the field $\mathbb{F}_{4}$, when considered under addition, is isomorphic to $\mathbb{F}_{2}\times\mathbb{F}_{2}$. Exploiting this structure, Ref. [BLOKHUIS2004161] demonstrates that the check matrix $H$ of an additive code with parameters $(\chi_{I},2^{k},d)_{4}$ can be constructed from a collection of $\chi_{I}$ lines in the projective space $\rm PG(2\chi_{I}-k-1,2)$, such that any d1d-1 lines in the set are linearly independent.
To universally decouple two-local interactions, we require the code $\mathcal{C}$ to have distance $d=3$. Consequently, we need to choose a collection of $\chi_{I}$ lines in the projective space $\rm PG(2\chi_{I}-k-1,2)$ such that any $2$ lines in the set are linearly independent. This set is known in projective geometry as a (partial) 1-spread. The maximum logical dimension $k$ is given by
@@
where $h$ is the smallest integer number such that $2^{h}\geq 3\chi_{I}+1$ for even $h$, and $2^{h}\geq 3\chi_{I}+5$ for odd $h$. As a result, the orthogonal array constructed from $\mathcal{C}^{\perp}$ with $H$ as the additive generator matrix has parameter
[…]
As a simple example, Table 2(a) lists a full 1-spread of the projective space $\rm PG(3,2)$, where each projective line is represented by two projective points. Each line is further associated with a color class $c_{i}$. In Table 2(b), we map each projective line in $\rm PG(3,2)$ to a projective point in $\rm PG(3,4)$ via a simple embedding from $\mathbb{F}_{2}\times\mathbb{F}_{2}$ to $\mathbb{F}_{4}$
@@
The resulting rows in the table form generators of a decoupling group capable of suppressing arbitrary two-local interactions.
We note that the above construction uses the same number of color classes and length in the decoupling sequence as the multi-axis protocol introduced in Ref. [brown2024efficient]. This is not a coincidence. The multi-axis protocol can be understood as implicitly employing an additive code over $\mathbb{F}_{4}$. This connection arises because the Hadamard matrix defines a binary linear code (the Hadamard code), whose rows or columns correspond to codewords. By grouping these rows into Schur sets [leung2002simulation] and mapping from $\mathbb{F}_{2}\times\mathbb{F}_{2}$ to $\mathbb{F}_{4}$, one preserves the additive structure of the binary code. Moreover, Ref. [votto2024universal] shows in their work (under the name of Walsh functions) that with minor modifications, the same construction can be extended to enable universal quantum simulation. Consequently, our framework naturally includes the protocols from Ref. [brown2024efficient, votto2024universal], providing an alternative way to understand their underlying structure.
@@
Figure 7: Scaling of decoupling sequences. Scaling of sequence length $L$ with the chromatic number $\chi_{I}$ for the code constructions listed in Table 3. (a) For comparison, the linear baseline $L=4\chi_{I}$ is shown in black. All codes with distance $d=3$ exhibit linear scaling in $\chi_{I}$, while those with $d=4$ show quadratic scaling, consistent with the scaling reported in Ref. [Bookatz_2016].
@@ V.2.3 Tailored sequences for spin-qubit
A similar strategy also applies to additive projective geometry codes: for each line in the projective space, there exist three distinct representations that also avoid Heisenberg exchange terms as codewords. As an illustration, Table 2(c) presents a tailored DD group designed to suppress Heisenberg exchange interactions, derived from the construction in Table 2(b). Notably, this construction increases the number of color classes from 5 to 15 while retaining the same number of generators. In the case of the multi-axis protocol introduced in Ref. [brown2024efficient], this modification is equivalent to cyclically permuting the rows in the Schur set.
@@
Let us apply this idea to the quotient hypergraph in Fig. 6(b). We begin by constructing a cap set in $\rm PG(2,4)$, which corresponds to a universal DD sequence that decouples all three-local interactions among six qubits. This cap set is given by the additive generators of the self-dual hexacode [conway2013sphere]
@@
The resulting DD sequence suppresses both the Heisenberg exchange interactions and the scalar chirality terms. However, it stills contains six generators. By mapping the generator set to a minimum covering problem, as described in Sec. IV.4, we find that only four of these generators are necessary to suppress all relevant interaction terms
@@
We emphasize that the expansion approach is effective primarily for bang-bang sequences. Under bounded control, finite pulse durations can lead to Heisenberg exchange or scalar chirality interactions transforming into other three-local terms that the sequence may not suppress. In such cases, alternative constructions may be required to obtain tailored, time-efficient bounded-control DD sequences. Alternatively, one may resort to the universal DD sequences discussed in Sec. V.2.1, which, although potentially longer, suppress all three-local interactions. Finally, we note that shorter sequences can often be obtained by systematically removing generators from a universal sequence, one at a time, as demonstrated above. This process might reduces the number of supported color classes but preserving the required decoupling properties. Developing such compact, tailored sequences remains an interesting direction for future work.
@@
We also analyze the sequence lengths as a function of the chromatic number $\chi_{I}$. To visualize this, in Fig. 7(a) we show the scaling of different DD sequences that suppress two-local interactions as the chromatic number $\chi_{I}$ increases; in Fig. 7(b), we focus on sequences targeting three-local interactions. In both cases, we compare the scaling behavior of the sequence lengths against ideal linear and quadratic scaling. For two-local interactions, we find that the universal constructions scale linearly with sequence length $4\chi_{I}$, while the tailored constructions scale more compactly with $2\chi_{I}$. In the case of three-local interactions, both universal and tailored constructions exhibit quadratic scaling $L\propto\chi_{I}^{2}$, but the tailored sequences have a smaller prefactor, leading to more efficient implementations. These scaling behaviors are consistent with those of Ref. [Bookatz_2016].
@@ V.3 Digital-analog simulation
In the previous section, we showed how to construct DD sequences that suppress residual interactions within the system and its environment. In this section, we demonstrate how such sequences can also be leveraged for Hamiltonian simulation and digital-analog quantum computing. For concreteness, we consider the Kitaev’s honeycomb lattice toy model, described by the Hamiltonian [kitaev2006anyons]
@@
where the sum is taken over the red (r), blue (b), and green (g) colored edges in Fig. 8(a). This model is exactly solvable by using a mapping to Majorana fermions and exhibits both gapped and gapless phases depending on whether the coupling strengths $J_{x},J_{y},J_{z}$ satisfy the triangle inequality [kitaev2006anyons]. Here, we focus on simulating the system deep in the gapless phase, where $J_{x}=J_{y}=J_{z}=J$.
@@
Figure 8: Dynamical quantum simulation of the Kitaev model. (a) A finite-size Kitaev honeycomb lattice, where edges are colored red, green, and blue to represent $X_{i}X_{j}$, $Y_{i}Y_{j}$, and $Z_{i}Z_{j}$ interactions, respectively. The labels 11 through 66 indicate the equivalence classes of qubits under the vertex color partition which respects the edges coloring, not individual physical qubits. (b) The folded Kitaev lattice (i.e., the quotient hypergraph) obtained by collapsing all qubits within the same equivalence class. Compared to the standard unit cell, the quotient hypergraph includes three additional edges (2-5), (3-4), and (1-6) that arise from the folding.
As in previous sections, we begin by identifying the effective number of equivalence color classes in Kitaev’s honeycomb model. If we were to universally decouple all interactions, the honeycomb lattice would be trivially 2-colorable. However, because we aim to selectively preserve certain interactions, the effective model becomes more complex.
In particular, due to the link-dependent interaction terms in the Hamiltonian, the symmetry of the model reduces the problem to an effective system involving six color classes. In Fig. 8(a), we color the qubits that are equivalent under the translation symmetry of the honeycomb lattice. The resulting quotient hypergraph, formed by identifying these symmetry-equivalent sites, contains six color classes. There are three extra edges appear between opposite effective vertices, as shown in Fig. 8(b).
To demonstrate the power of the effective graph in Fig. 8(b), we use the simple algorithm presented in Sec. IV.4 to find the minimal decoupling group for the bang-bang sequence. We also explicitly construct a decoupling group for the bounded sequence, see Appendix D. Using the same terminology as before, we associate $H_{S}^{\parallel}=H_{\rm comb}$. Consequently, the remaining interactions belong to the list $H_{S}^{\perp}$.
@@
$W_{3}=Y_{1}Y_{3}X_{5}X_{6},\quad W_{4}=Y_{2}X_{3}X_{4}Y_{5},$ (57b)
and are closely related to a known symmetry of the Kitaev honeycomb model, namely the plaquette operators [kitaev2006anyons]
$W_{p,1}=Z_{1}X_{2}Y_{3}Y_{4}X_{5}Z_{6},~\quad W_{p,2}=Y_{1}Y_{2}Z_{3}X_{4}Z_{5}X_{6}$ (58)
One can verify that $W_{1}W_{2}=W_{p,1}$ and $W_{3}W_{4}=W_{p,2}$, that is, each pair provides a decomposition of the original plaquette symmetry. The emergence of these subsymmetries is expected: since the effective Kitaev honeycomb model arises from folding the original lattice, it can exhibit additional symmetries not present in the unfolded geometry.
If we take $W_{p,1}$ as the generator of $\mathcal{G}$, the resulting sequence effectively removes the interaction terms along the outer edges of the hexagon in Fig. 8(b), aligning the system’s dynamics with that of the Kitaev model. However, the inner-edge interactions commute with $W_{p},1$, so they are not eliminated by this sequence and remain in the effective Hamiltonian. This again reflects the presence of the subsymmetries introduced by lattice folding. In a similar manner, if we take $W_{p,2}$ as the generator of $\mathcal{G}$, there are interactions on outer edges that commute with $W_{p,2}$.
Either by numerically solving the associated minimum-set covering problem or by manually verifying it, we find that using either the pair $(W_{1},W_{2})$ or $(W_{3},W_{4})$ as generators is sufficient to construct a DD group $\mathcal{G}$ for the bang-bang sequence. Consequently, the effective Hamiltonian have the same interaction as shown in the edges of Fig. 8(b). However, the global site is opposite of what it should be, i.e $H_{S}^{\parallel}=-H_{\rm comb}$.